### v1

# Measurement of ion acceleration and diffusion in a laser-driven magnetized plasma

Abstract
Here we present results from an experiment performed at the GSI Helmholtz Centre for Heavy Ion Research. A mono-energetic beam of chromium ions with initial energies of MeV was fired through a magnetized interaction region formed by the collision of two counter-propagating laser-ablated plasma jets. While laser interferometry revealed the absence of strong fluid-scale turbulence, acceleration and diffusion of the beam ions was driven by wave-particle interactions. A possible mechanism is particle acceleration by electrostatic, short scale length kinetic turbulence, such as the lower-hybrid drift instability.
Cosmic rays (CRs) and their extraterrestrial origin were discovered more than a century ago [1], with measurements of the CR energy spectrum up to eV [2, 3]. The exact mechanism that leads to such high energy particles remains controversial, as does the nature of the sites where this acceleration occurs. Although many different processes may result in CR acceleration, the present day understanding is that turbulence plays an essential role in energizing both electrons and ions. The original mechanism of CR acceleration by turbulence or plasma waves was proposed by Fermi [4], where CRs are produced as charged particles gain energy in random scattering events with magnetized clouds. Like turbulence, Fermi acceleration is expected to be ubiquitous in the universe, but laboratory evidence for the process remains elusive.
In addition, inhomogeneous astrophysical environments are susceptible to various instabilities, such as the lower-hybrid drift instability (LHDI) [5] which arises from plasma gradients. These instabilities can accelerate particles, possibly playing an important role in the overall energizing of the thermal background [7, 6, 8, 9]. Space probes have detected lower-hybrid waves (LHWs) in the terrestrial magnetotail current sheet and their role in the energy dissipation of electrons [10]. A wave-particle model involving LHWs has been used to explain the presence of accelerated auroral electrons [11] and ions [12], and similarly for the simultaneous acceleration of electrons and ions in solar flares [13, 14].
Laboratory astrophysics experiments are an invaluable tool in studying complex phenomena such as these acceleration mechanisms [15, 16, 17], used in complement with simulations while addressing some of their limitations. Compared to in situ measurements, where many competing processes can obfuscate the physics that is happening, laboratory plasmas can be designed to study specific mechanisms of interest. This paper presents experimental results from an investigation on ion acceleration and diffusion in a magnetized laser-driven plasma conducted at the GSI Helmholtz Centre for Heavy Ion Research. The platform was a modified version of that previously fielded to study the turbulent dynamo [18, 19, 20, 21, 22, 23], coupled with the heavy ion accelerator at GSI. Plasma conditions were measured and supplemented by results from radiation-magneto-hydrodynamic (MHD) FLASH simulations [24].
Experimental setup— The experimental platform employed for the experiment at the GSI Helmholtz Institute for Heavy Ion research is shown schematically in Fig. 1. Supersonic plasma jets are produced via laser ablation of two opposing target foils, each machined with an embedded grid structure to impose density perturbations. Simultaneously, seed magnetic fields in the azimuthal direction are generated by the Biermann battery mechanism from the cross-product of the density and temperature gradients [25, 26, 27]. The two jets collide at the midpoint between the two foils, defined as the target chamber center (TCC), forming an approximately cylindrical magnetized interaction region. The intention of the design was for the initial density perturbations advected by the jets to interpenetrate, introducing shearing motions which would drive turbulent mixing of the plasma. The scale of these instabilities is however, predicted to be smaller than the experimental resolution of the diagnostics [24]. Additionally, by only driving a single target foil, the case of reduced fluctuations, plasma density, and field strength was accessed for comparison. Interferometry and ion deflectrometry were fielded as diagnostics for the experimental plasma conditions.
The monoenergetic ion beam from the UNIversal Linear ACcelerator (UNILAC) was directed through the interaction region, acting as a surrogate for CR particles. The UNILAC supplied a Gaussian pulse train (4 ns width, 28 ns spacing), with only one pulse traversing the interaction region in each shot. The energy and initial charge state of the ion beam were tunable parameters and we present results obtained using the following: 52Cr14+ with energies of 8.75 MeV/u; 52Cr20+ with energies of 8.661 MeV/u.
The ion beam had a diameter of mm at the TCC, with a divergence of mrad. A section of the accelerator beamline, located downstream of the laser interaction chamber, was used to direct the ions toward a diamond detector [28, 29] situated approximately 12 m beyond the interaction region. This arrangement enabled the acquisition of ion time-of-flight (ToF) data.
Results— Line-integrated electron densities in the ablated plasma flows and interaction region were measured using laser interferometry, providing a coarse measurement of the plasma density structure. This data also provided an upper bound for the level of fluid turbulence by considering finer density fluctuations below the smaller scale width of the plasma ( mm). In addition, the density profile of the plasma expansion was used to estimate the temperature.
The fielded setup consisted of an imaging Mach–Zehnder interferometer and a 355 nm, 500 mJ, 500 ps pulsed probe laser. This data was time-gated by the laser-pulse duration, which is short compared to the system’s dynamical timescale ( ns), where is the sound speed and mm is the observed experimental system size, equivalent to the path lengths of the ion beam and interferometry probe through the interaction region. The evolution of the plasma can be tracked with interferometry by changing the probing time in separate shots.
The free electron density, , is retrieved from the raw interferograms via the following process. Firstly, the fringes were traced manually to form phase contour maps. In regions associated with strong density gradients, for example adjacent to the target foils, individual fringes cannot be resolved – these regions are excluded from the interpolation.
Next, the phase contour maps are interpolated to recover a map of the phase shift at each point [30, 31]. The phase shift is converted into the line-integrated electron density via the well known relation:
where is the number of fringe shifts, is the critical density of the plasma, and is the vacuum wavelength of the probing laser. The result is a 2D map of line-integrated electron density (Fig. 2C). At ns, a clear interaction region has formed with a width of mm. Given the approximate azimuthal symmetry of the setup, the path length traversed by the probe beam is also 2 mm, resulting in an average volumetric electron density in the interaction region of cm-3. Assuming full ionization with the given plasma conditions, consistent with atomic kinetic simulations, the corresponding ion density cm-3. Similar analysis for data recorded at ns gives values of cm-3 and cm-3.
The evolution of the inflow plasma jets was captured by interferometry on shots with only one target foil driven by nhelix. The expanding jet can be approximated as isothermal, since the ratio of the temperature equilibration timescale to the experiment duration can be shown to be [32] ; equivalently this implies a small Péclet number (). An isothermal expansion has a known density profile [33, 34] dependent on the plasma sound speed . Since is related to the plasma temperature [35], at a time of ns, approximately the time the jets collide, the temperature of the inflow jets from this analysis is eV. Even accounting for collisional heating, we can expect the interaction region plasma to be in the temperature regime of the low hundreds of eV.
Qualitatively, the interferometry data in Fig. 2 shows a relatively laminar interaction region, with no evidence of large-scale density fluctuations associated with strong fluid turbulence. To bound any residual fluctuation, we low-pass filter each traced fringe at a cutoff set by the width of the interaction region, take the result as a slowly varying baseline, and subtract it from the raw trace. The residual displacement is then mapped via Eq. (1) to the corresponding fluctuation in line-integrated electron density.
Next, assuming an isotropic plasma, it can be shown [19] with a random-walk treatment that the line-integrated density fluctuation is related to the volumetric density fluctuation by:
where is the characteristic length scale of the density fluctuations (approximated as the grid spacing [19]) and is the path length of the interferometry laser through the plasma. The magnitude of can be used to infer the turbulent velocity using a continuity argument in subsonic, incompressible flow [36]:
Using the values of predicted by MHD simulations [24] for the case of colliding plasma flows, which are more than twice those extracted from the single-sided experimental fits, this results in values of km s-1 and km s-1 at and 15 ns, respectively. These values should be interpreted as an upper bound for , as the effect of limited dynamic range as well as the variability of the manual tracing process cannot be isolated from actual density fluctuations. This analysis will in turn provide upper bounds for the expected magnitude of second order Fermi acceleration, shown in later discussion.
Another important parameter relevant to transport properties in the plasma is the magnetic field strength, which was constrained experimentally by fielding an ion deflectrometry diagnostic. A nuclear-track detector (CR-39) was placed in the path of the ion beam cm after the TCC. This was fielded in lieu of the ToF diamond detector and data were collected both with and without the laser-ablated plasma. For these measurements, the ion beam was gated to the minimum possible length of 10 µs, approximately 350 pulses, to minimize the number of undeflected pulses incident on the plate.
Spatially resolved ion fluence was recorded with this setup. The distribution of pits demarcating the position of ion impacts consists of a saturated central region containing the bulk of the ion beam and a surrounding ‘halo’ decreasing in pit density with increasing radial distance; Fig. 4A and 4B show a section of the data aligned radially outwards from a knife edge positioned mm from the center of the ion beam, for shots without and with plasma respectively. As the ion beam traverses the plasma, it accumulates a random-walk deflection proportional to the magnetic field strength, which dominates over contributions from the electric fields and Coulomb collisions. The relative contributions from magnetic and electric fields was compared by considering the Lorentz equation of motion and applying resistive MHD; the ratio of deflection force from electric and magnetic fields , where is the characteristic fluid velocity and the initial ion beam velocity.
Since the beam ion gyroradius m is much larger than the magnetic coherence length m, the ions are unmagnetized and their deflection accumulates in a series of small-angle scattering events as a random walk process; this manifests as a Gaussian deflection as a consequence of the central limit theorem. The average angle of deflection is [37]
where is the total number of scattering events and can be approximated as . The RMS deflection angle is thus [38]
where , and are the charge state, mass and energy of the beam ions. An upper estimate for the magnetic coherence length is found by taking the grid spacing (60 um) on the target foils and accounting for the approximate doubling in spatial extent of the plasma: mm. This is on the order of the outer scale of turbulence, which in previous experiments (on similar setups) has been larger than [19]. Using the experimental value of obtained from the width of the distribution in Fig. 4D, this gives a minimum field strength of kG. An upper bound for the field strength can be defined using a simple geometric argument: assuming all the counts in the distribution of Fig. 4D come from ions at the center of the beam, we obtain which is found by summing the initial ion beam divergence and the measured . This corresponds to a field strength of kG. Thus, we estimate kG, which has implications for later discussion on acceleration from fluid and short-scale electrostatic turbulence.
| Drive | Single | Double | |||
| (ns) | 5 | 10 | 15 | 10 | 15 |
| (cm-3) | |||||
| (cm-3) | |||||
| (eV) | — | — | |||
| (kG) | — | — | — | 40-230 | — |
| (cm s-1) | |||||
| (cm s-1) | — | — | |||
| (cm-3) | — | — | |||
| (cm-3) | — | — | |||
| (eV) | 412 | — | — | 106 | 80 |
| (kG) | 38 | — | — | 77 | 40 |
| (cm s-1) | — | — | |||
1
FLASH simulation data was retrieved with the following method. Values are averaged over a 0.95 mm radius cylinder centered between the targets, which extends the extent of the domain. A temperature threshold is used to exclude solid target material. Values are averaged over two runs, with 25J and 35J of laser energy on each target foil, respectively.
The ToF diamond detector fielded allows for an energy measurement of each pulse within the ion beam. Figures 5A and 5B present the resulting energy spectra, comparing the pulse of interest (PoI), that is, the ion pulse that coincides with the plasma, with a background averaged over pulses acquired prior to plasma formation. There is significant shot-to-shot variation even for shots with similar plasma conditions probed at the same time, reflecting the inherently stochastic nature of the beam-plasma interaction. To quantitatively compare shots across the campaign, two main metrics were used: the shift in mean energy of the pulse , calculated by integrating the energy profile to find its center-of-mass; and the change in the energy spread of the pulse, defined as twice the standard deviation of the pulse around its mean energy.
Fig. 6 shows the distribution of the measured acceleration and diffusion from this experiment, revealing a clear distinction between single- and double-sided drive configurations. Single-sided shots cluster near the origin, exhibiting minimal energy shift and broadening, whereas double-sided shots trace a broader curve with significantly larger magnitudes in both metrics. Nevertheless, outliers are present in both cases, reflecting the inherent stochasticity of the beam–plasma interaction. Notably, the double-sided data appears to be divided into two regimes: one characterized by net acceleration with minimal broadening, and another by increased width accompanied by a reduction in mean energy.
In addition, charge transfer processes between the beam and plasma ions introduces a non-trivial correction to as the ion beam traverses the plasma. To account for this effect, a semi-empirical model [39] based on the Gus’kov model [40] was employed to estimate the mean non-equilibrium charge state of the beam ions. This upshifted value is accounted for in the estimate (see Table I).
One possible explanation for deceleration is the charged-particle stopping power of the plasma — the magnitude of this effect can be calculated following a generalized Fokker-Planck equation [41, 42] which accounts for both small and large-angle binary collisions:
where is the plasma frequency and is the Coulomb logarithm. Using the experimentally constrained plasma conditions in Table I, the electron-ion stopping dominates, with values of and MeV for and , respectively, assuming a path length of 2 mm. This magnitude of deceleration cannot account for the observed shifts in mean energy. Similarly, the expected parallel diffusion associated with Coulomb collisions is found to be negligible ( eV) under the experimental conditions, due to the high initial velocity of the ion beam.
The results can be compared with theoretical predictions, considering two candidate mechanisms informed by previous theoretical work [24]. The first is second-order Fermi acceleration [43], originally proposed to account for the origin of high-energy cosmic rays by interactions between particles and mirror fields in the interstellar medium [4]. This process can be generalized to a diffusive interaction between charged particles and magnetized plasma turbulence. For this experiment, as the ion beam traverses the interaction region, the ions undergo many deflections by magnetic fluctuations which result in a total energy change of
where is the speed of light and and are the RMS velocity and magnetic field, respectively. Taking (bounded by deflectrometry) and (bounded by interferometry), the maximum associated with second-order Fermi acceleration (see Table I) is much smaller than the initial energy spread of the ion beam ( MeV) and the energy resolution of the diamond detector. Consequently, while this mechanism may contribute, it cannot be responsible for the changes measured.
The second possibility is a wave-particle interaction, such as the lower-hybrid drift instability. LHDI is a kinetic instability driven by strong density and magnetic field gradients [5], which can excite lower-hybrid waves (LHWs) that in turn induce an electric field for ion acceleration [6]. To demonstrate that the experimental data is not inconsistent with the formation of the instability, we infer an upper bound for the LHDI growth rate using experimental values. The maximum linear growth rate of this instability is given by
where mm is the gyroradius of the plasma ions, is the lower-hybrid frequency, and is the inverse gradient scale length which can be approximated as
Using the magnitude of density fluctuations bounded by interferometry, s-1 at ns, corresponding to an e-folding time of 1.6 ns. Additionally, the gradient scale length mm satisfies the steepness condition [44] required to excite LHDI. Note that this analysis only includes density gradients associated with resolvable fluctuations, limited by the dynamic range of the interferometry diagnostic. Gradients on a smaller scale closer to would instead excite kinetic instabilities, with higher growth rates on the order of rad s-1 [44, 45, 46].
In the context of this experiment, LHWs are assumed to interact linearly and non-resonantly with the ion beam in a weakly turbulent plasma [47]. This can be treated as a diffusion problem, with an associated energy change of
where C ms is a constant of proportionality that captures the level of lower-hybrid turbulence (see [24] for details) and u the atomic mass of the beam ions. The theoretically predicted values (Table I) for LHDI are sufficiently high for measurement and are expected to dominate over the effect of second-order Fermi acceleration.
The observation of bulk energisation cannot be explained by the broadening process outlined above alone, and suggests that coherent fields which facilitate particle acceleration are generated in the plasma. A tantalising possibility consistent with our experiment is the role of short scale electrostatic turbulence (below the interferometry diagnostic’s resolution) in developing coherent structures within the plasma. In literature, both Langmuir wave [48] and LHW [49, 50] turbulence have been shown to drive the evolution of these wave structures.
In summary, we present observations of ion acceleration and diffusion in a laser-driven magnetized plasma. Despite the absence of large-scale fluid turbulence as diagnosed by interferometry, evidence for ion acceleration and energy diffusion was observed with a time-of-flight detector, suggesting that a wave-particle interaction was the underlying acceleration mechanism. Analysis showed that the lower-hybrid drift instability was consistent with the measured acceleration and diffusion, although other wave-particle interactions could also be responsible.